A tank was 1/2 full of water.after 420 gaallons had been taken out it was 1/8 full.Find the capacity of the tank.
1
Expert's answer
2012-04-03T08:28:45-0400
Let's denote capacity of tank as x. So we have:
1/2*x-420=1/8*x
(1/2-1/8)x=420
(4/8-1/8)x=420
3/8*x=420
x=420*8/3=340*8=1120 gallons